How do you find the sum of 7+10+13+..........+100 ???
Answer by Alan3354(69443)

How do you find the sum of 7+10+13+..........+100 ???
----------
You can add them.
----
Or, find the # of elements:
(100 - 7)/3 = 31 --> 32 elements
Make 16 pairs,
7 + 100 = 107
10 + 97 = 107
etc
16*107 = 1712
Uses less paper (and time) than adding them.